OpenStudy (anonymous):

Evaluate the expression below when x = -2.

OpenStudy (cggurumanjunath):

where is the expression ?

OpenStudy (anonymous):

expression?

OpenStudy (cggurumanjunath):

substitute x = -2 in your given expression !

OpenStudy (anonymous):

hold on Evaluate the expression below when x = -2. x2 - 8 x + 6 Evaluate the expression below when x = -2. x2 - 8 Evaluate the expression below when x = -2. x2 - 8 x + 6 -3 -1 1 2 3 2

OpenStudy (cggurumanjunath):

X ^ 2 -8 WHEN X=-2 ( -2 ) ^ 2 -8 = 4 -8 = -4

OpenStudy (cggurumanjunath):

SIMILARLY DO THE REMAINING.